A fruit juice company mixes pineapple juice that sells for $5.50 per gallon with 100 gallons of orange juice that sells for $3.00 per gallon. How much pineapple juice is used to make a pineapple-orange drink selling for $3.50 per gallon?
My answer is
$5.50 x + $3.00 (100) = $3.50 (100 +x)
$5.50x + $300 = $3.50(100 + x)
$5.50x + $300 = $350 +$3.50 x
$5.50x =$50 +$3.50x
2 x = 50
X=25. But the book says 75. Please help!

Answer by mananth(16949)

------ price ---------------- quantity
YOU ARE RIGHT

orange juice 3 ---------------- 100 gallons
pine apple juice 5.5 ---------------- x gallons
Total 3.5 ---------------- 100 + x gallons

3*100+5.5*x=3.5 (100+x)
300+5.5x =350+3.5 x

5.5 x - 3.5 x = 350 - 300

2 x = 50

/ 2

x = 25 gallons pine apple juice